OpenStudy (anonymous):

Need Help with this question please Tan^2x/secx=secx - cosx

OpenStudy (anonymous):

Can someone help me out please

OpenStudy (anonymous):

@Benny43 i will help

OpenStudy (anonymous):

(tan x)^2/sec x={(sec ^2 x) -1}/sec x=sec x-cos x.@Benny43

OpenStudy (anonymous):

@Benny43

OpenStudy (solomonzelman):

tan^2x/secx=secx - cosx tan^2x=sec^2x - 1

OpenStudy (solomonzelman):

I multiplied both sides times the secant

OpenStudy (solomonzelman):

and this tan^2x=sec^2x - 1 or tan^2x=sec^2x - 1

OpenStudy (solomonzelman):

is an identity, although we can prove that one too, if you want/need.